OpenStudy (anonymous):

An electron has a total energy that is 250 times its rest energy. determine its momentum

OpenStudy (matt101):

You can solve this equation using the energy-momentum relation: \[E^2_{total}=(mc^2)^2+(pc)^2\] The question tells us that E(total)=250mc^2, so we can plug this into the above equation to get: \[(250mc^2)^2=(mc^2)^2+(pc)^2\] The mass of the electron and speed of light are both constants that you probably have, so just plug in the values and solve for p!

OpenStudy (anonymous):

can i use the equation "p=gamma*m*u"?

OpenStudy (matt101):

As a matter of fact yes, I believe you would need to consider gamma!

OpenStudy (anonymous):

i have a question about the other problem, i dont know how to explain the equation" beta=(pc)/E" is a useful relation to find the velocity of a highly energetic particle.